this comparison is being made for a state government that mandates the use of benefit/cost criteria at an interest rate of 8%. Which ratio(s) should be calculated, and why? Based on your calculations, which mutually exclusive alternative should be chosen? Project B can be repeated in 5 years.

Alternative A B
First cost $200,000 $225,000
Annual O&M $25,000 $15,000
Salvage Value $0 $50,000
Life 10 years 5 years
Annual User Benefit $75,000 $85,000
Answer #1

Benefit-cost ratio, should be calculated, because it considers the benefits and cost associated with each alternative ans it is the criteria required by the decision makers.

For alternative A:

Uniform first annual cost = 200000*(A/P, 8%, 10) = 200000*.1490

Uniform first annual cost = $29800

Benefit-cost ratio = 75000/(29800+25000)

Benefit-cost ratio = 1.37

----

For alternative B:

Uniform first annual cost = 225000*(A/P, 8%, 5) = 225000*.2505

Uniform first annual cost = 56362.5

annual uniform salvage value = 50000*(A/F, 8%, 5) = 50000*.1705

annual uniform salvage value = $8522.82

Benefit-cost ratio = 85000/(56362.5 + 15000 - 8522.82)

Benefit-cost ratio = 1.35

Since Benefit-cost ratio for alternative A is higher than that of alternative B, so alternative A should be selected.